OpenStudy (anonymous):

/_\ DEF has vertices D(-2,4), E(0,-2), and F(6,2). Determine: a) an equation for DG, the median from D to EF b) an equation for FC, the altitude from F to DE c) an equation for AB, the right bisector of DF

jimthompson5910 (jim_thompson5910):

point G is the midpoint of segment EF so you use the midpoint formula to find point G

jimthompson5910 (jim_thompson5910):

what do you get when you do that

OpenStudy (anonymous):

Point G is (3,0)

OpenStudy (anonymous):

Then i found the slope of the of DG which is -4/5

jimthompson5910 (jim_thompson5910):

so far, so good

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

Im not sure

jimthompson5910 (jim_thompson5910):

well you know the slope -4/5

jimthompson5910 (jim_thompson5910):

you know a point that this line goes through (-2,4)

OpenStudy (anonymous):

the final answer is 4x+5y-12=0

jimthompson5910 (jim_thompson5910):

so y = mx + b y = (-4/5)x + b 4 = (-4/5)(-2) + b 4 = 8/5 + b 4 - 8/5 = b b = 12/5

OpenStudy (anonymous):

so do i put it into y=mx+b?

jimthompson5910 (jim_thompson5910):

the equation in slope intercept form is y = (-4/5)x + 12/5

jimthompson5910 (jim_thompson5910):

if you want the equation in standard form, you would get y = (-4/5)x + 12/5 5y = -4x + 12 4x + 5y = 12

OpenStudy (anonymous):

thanks alot!!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome